I am struggling with these types of problems
x/5+x=-2
do you know an easier way to do these

Answer by checkley71(8403)

X/5+X=-2
(X+5X)/5=-2
X+5X=-10
6X=-10
X=-10/6 OR X=-5/3
PROOF
(-5/3)/5-5/3=-2
(-5/3*1/5)-5/3=-2
-5/15-5/3=-2
-1/3-5/3=-2
-6/3=-2
-2=-2